# Cog-GA: A Large Language Models-based Generative Agent for Vision-Language Navigation in Continuous Environments

### v1

@@ II Related work
II-A Vision-Language Navigation
Visual language navigation (VLN) is an emerging interdisciplinary field that has garnered significant attention from natural language processing, computer vision, and robotics communities. Anderson et al. first proposed the VLN task in 2018, introducing the Room-to-Room (R2R) dataset based on real-world environments [2]. Following the release of the R2R dataset, various expansions of VLN tasks emerged, such as the outdoor visual language navigation Touchdown [5] and Remote Embodied Visual Referring Expression in Real Indoor Environments (REVERIE) [18]. Numerous methods have been proposed to address these tasks. A notable advancement is the Reinforced Cross-Modal Matching (RCM) approach, which has outperformed baseline methods by 10%
The History Aware Multimodal Transformer (HAMT) has set a new benchmark in long-term navigation, demonstrating the importance of incorporating historical context in navigation tasks [7].
II-B VLN in Continuous Environments
@@
In traditional VLN tasks, agents navigate through a restricted graph, an unrealistic assumption for real-world navigation robots. Krantz et al. [15] extended the discrete R2R VLN task setup to continuous environments, where agents make navigation decisions in freely traversable 3D spaces. The VLN-CE framework introduces new challenges and brings the task closer to real-world navigation scenarios. Zhang et al. [27] highlighted the significant impact of visual appearance features on agent performance, underscoring the need for models that generalize better across diverse environments. Similarly, Wang et al. [25] introduced the Reinforced Cross-Modal Matching (RCM) approach, achieving notable improvements in navigation performance. Guhur et al. [9] demonstrated that in-domain pretraining using the BnB1 dataset significantly enhances generalization to unseen environments.
@@
To address the complexities of the VLN-CE task, Hong et al. [10] showed that agents navigating with predicted waypoints perform significantly better, reducing the discrete-to-continuous gap. Wang et al. [24] proposed an Environment Representation Graph (ERG) that strengthens the relationship between language and environment, leading to improved performance in VLN-CE tasks. Chen et al. [6] introduced DNA, a direction-guided navigator agent that integrates directional cues from instructions into the encoder-decoder framework. However, current VLN-CE methods require massive training to obtain prior knowledge.
II-C Large Language Models Guided Navigation
Leveraging their powerful information processing capabilities and extensive prior knowledge, large language models (LLMs) have emerged as a novel approach for reasoning in navigation tasks. The VELMA model [21] uses LLMs to predict movement directions based on landmark information from language instructions. The Esc model [28] considers correlations with target objects at both the object and room levels. For zero-shot object navigation, LFG [22] introduced chain-of-thought (CoT) reasoning in LLMs to avoid navigating to irrelevant areas. Cai et al. [3] extended the use of CoT by clustering panoramic images into scene nodes and employing CoT to decide whether to explore or exploit, selecting images most likely to contain the target object and navigating accordingly. The Prompt-based Environmental Self-exploration (ProbES) [16] significantly enhances LLMs’ generalization capabilities in tasks like VLN and REVERIE, indicating a shift towards more adaptive and context-aware LLMs in navigation. Co-NavGPT [26] applies LLMs’ decision-making processes in multi-robot collaborative navigation, centrally planning midterm goals for each robot based on online map information. However, more cognitive processes in the brain can also be introduced to this process.
@@ III-A Generative Agent for VLN-CE Tasks
We divide VLN-CE tasks into three phases: generating the search space, high-level target planning, and low-level motion generation. Constructing the search space, a crucial preprocessing step, segments the continuous environment into waypoints, simplifying navigation to point selection and improving efficiency. We use a planner based on large language models (LLMs) for high-level target planning. It selects a waypoint as the next target based on current sub-instructions using spatial memories from the cognitive map of the memory stream. The target waypoint is then sent to the motion generator for action execution.
We propose a generative agent for VLN-CE tasks, including the waypoint predictor, memory stream, instruction processing, high-level planner, and reflection module. The instruction processing module divides the task into shorter-range sub-instructions. The waypoint predictor uses panoramic observations to construct the waypoint search space at each step. A scene describer characterizes the observed scene, dividing descriptions into the ”what” stream (landmark objects) and the ”where” stream (spatial characteristics). These descriptions adjust the sub-instruction for alignment with the environment, allowing the planner to focus on the current target: cognitive maps and reflection memories from the memory stream guide the high-level planner. The planner uses scene descriptions, memories, and the sub-instruction to compose a prompt for the LLM. The target waypoint index is conveyed to the low-level actuator. During movement execution, a reflection generator assesses the navigation results, constructing feedback to evaluate each navigation impact of step.
@@ III-B Cognitive Map based Target Inference
Retrieving the cognitive map is crucial for target inference. As shown in Figure 3, we define two retrieval methods: the history and observation chains. The history chain focuses primarily on navigated nodes, providing planners an abstract view of the current path. In contrast, the observation chain focuses on potential targets between the current and previous positions, offering a broader view of past decisions.
@@ III-D Generative Agent with Reflection Mechanism
After the planner identifies the target, the angle and distance to the waypoint are sent to the low-level actuator to move the agent. The agent then reflects on its movements to gather valuable experiences for future tasks. This reflection helps the agent understand why it succeeded or failed, gaining general knowledge about the environment. However, LLMs can be overwhelmed by too many experiences, disrupting their decision-making. The forgetting mechanism and evaluation metric for reflection memory are introduced by managing the stored information to address this issue. The agent calculates the score of each reflection memory. The bottom 10% of reflections will be eliminated.
@@ Acknowledgements
Appendix A Algorithm pseudocode
Appendix B Task setup
@@ 1. Task Setup
Appendix C Optimal Prompt Mechanism for LLMs in Navigation Tasks
@@ 2. Optimal Prompt Mechanism for LLMs in Navigation Tasks
Appendix D The Influence of Instruction Quality and Constructing Better Sub-Instructions
@@ 3. The Influence of Instruction Quality and Constructing Better Sub-Instructions
Appendix E The evaluation metric of reflection memory
@@ 4. The evaluation metric of reflection memory
Appendix F Vision-Language Navigation Task Sample

### v2

@@ II Related work
II-A VLN in Continuous Environments
Visual language navigation (VLN) has gained prominence across natural language processing, computer vision, and robotics. Introduced by Anderson et al. in 2018 with the Room-to-Room (R2R) dataset [2], VLN has since expanded to include tasks like Touchdown [5] and REVERIE [20] in diverse environments. Among the numerous methods developed, the Reinforced Cross-Modal Matching (RCM) method has notably surpassed baselines by 10% in the Success Rate weighted by Path Length metric [28]. The History Aware Multimodal Transformer (HAMT) further advances long-term navigation by effectively integrating historical context [7].
@@ II-A VLN in Continuous Environments
In traditional Vision-and-Language Navigation (VLN) tasks, agents are confined to a restricted graph, limiting their applicability to real-world scenarios. Krantz et al. [17] expanded the VLN paradigm to continuous environments, enabling agents to navigate through 3D spaces without pre-defined paths, as introduced in the VLN-CE framework. This setup more closely resembles real-world conditions but introduces new challenges. Zhang et al. [30] noted that visual appearance significantly affects agent performance, highlighting the necessity for models that generalize well in diverse settings. Wang et al. [27] developed the Reinforced Cross-Modal Matching (RCM) approach, which improved navigation performance. Additionally, Guhur et al. [9] showed that pretraining on the BnB1 dataset enhances model generalization to new environments.
The [13] model employs a dual-level decision-making process that uses high-level reasoning to match navigation instructions with visual cues and low-level policies for direct agent control. [12] introduced the Semantically-aware Spatio-temporal Reasoning Agent (SASRA), combining semantic mapping with a hybrid transformer-recurrence architecture to develop a temporal semantic memory, improving spatial and temporal reasoning capabilities in VLN.
@@
To further address the challenges of the VLN-CE task, Hong et al. [10] demonstrated that navigation via predicted waypoints markedly enhances performance by bridging the discrete-continuous divide. Wang et al. [26] proposed the Environment Representation Graph (ERG), which fortifies the connection between linguistic and environmental data, boosting VLN-CE performance. Chen et al. [6] introduced the Direction-guided Navigator Agent (DNA), integrating directional cues into the encoder-decoder framework. However, they noted that extensive training is necessary to acquire prior knowledge.
II-B Large Language Models Guided Navigation
Large language models (LLMs) enhance navigation tasks with robust information processing and extensive knowledge. The VELMA model [23] uses LLMs for navigation based on landmarks described in instructions. The Esc model [31] analyzes object and room-level correlations with targets. The LFG model [24] employs chain-of-thought (CoT) reasoning in LLMs for zero-shot object navigation, minimizing irrelevant travel. Cai et al. [3] extend CoT by clustering panoramic images into nodes for strategic navigation decisions. The Prompt-based Environmental Self-exploration (ProbES) [18] advances LLM generalization in VLN and REVERIE tasks, showcasing adaptability. Co-NavGPT [29] utilizes LLMs for collaborative multi-robot navigation, setting midterm goals based on live map data. Integrating additional cognitive processes could further improve these models.
@@ III-A Generative Agent for VLN-CE Tasks
We categorize the Vision-and-Language Navigation in Continuous Environments (VLN-CE) task into three phases: generating the search space, high-level target planning, and low-level motion generation. Initially, the search space is constructed by segmenting the continuous environment into waypoints, a crucial preprocessing step that simplifies navigation by reducing it to point selection, thereby enhancing efficiency. For high-level target planning, we employ a planner based on large language models (LLMs), which choose waypoints as targets based on current sub-instructions, utilizing spatial memories from the memory stream’s cognitive map. The selected waypoint is then forwarded to the motion generator for action execution.
We introduce a generative agent for VLN-CE tasks that comprises a waypoint predictor, memory stream, instruction processing module, high-level planner, and reflection module. The instruction processing module breaks down tasks into manageable sub-instructions, while the waypoint predictor constructs a search space for waypoints at each step using panoramic observations. A scene describer identifies and categorizes the environment into ”what” (landmark objects) and ”where” (spatial characteristics) streams, enhancing sub-instruction alignment with the environment. These streams and cognitive and reflection memories from the memory stream guide the high-level planner in forming prompts for the large language model (LLM). The planner uses these prompts to identify the target waypoint index relayed to the low-level actuator. During movement, a reflection generator evaluates the navigation results, providing feedback on each step’s impact.
@@ III-B Cognitive Map based Target Inference
Retrieving the cognitive map is crucial for target inference. As shown in Figure 3, we define two retrieval methods: the history and observation chains. The history chain focuses primarily on navigated nodes, providing planners with an abstract view of the current path. In contrast, the observation chain focuses on potential targets between the current and previous positions, offering a broader view of past decisions.
@@ III-D Generative Agent with Reflection Mechanism
After the planner identifies the target, the angle and distance to the waypoint are sent to the low-level actuator to move the agent. The agent then reflects on its movements to gather valuable experiences for future tasks. This reflection helps the agent understand why it succeeded or failed, gaining general knowledge about the environment. However, LLMs can be overwhelmed by too many experiences, disrupting their decision-making. We define three parameters for reflection memory in navigation tasks: optimal distance (measured by dynamic time warping between current and correct sequences), proximity (time closeness to the current step), and repeatability (frequency of similar memories). Identical new memories are not stored but update the existing memory’s proximity and repeatability. The score of each reflection memory is defined as follows:
@@ Acknowledgements
Appendix A Task setup
@@ 1. Task Setup
Appendix B Optimal Prompt Mechanism for LLMs in Navigation Tasks
@@ 2. Optimal Prompt Mechanism for LLMs in Navigation Tasks
Appendix C The Influence of Instruction Quality and Constructing Better Sub-Instructions
@@ 3. The Influence of Instruction Quality and Constructing Better Sub-Instructions
Appendix D The evaluation metric of reflection memory
@@ 4. The evaluation metric of reflection memory
Appendix E Vision-Language Navigation Task Sample